OpenStudy (anonymous):

The label on the car's antifreeze container claims to protect the car between −30°C and 130°C. To convert Celsius temperature to Fahrenheit temperature, the formula is C = 5 over 9 (F − 32). Write and solve the inequality to determine the Fahrenheit temperature range at which this antifreeze protects the car. −30 > 5 over 9 (F − 32) > 130; −22 > F > 266 −30 < 5 over 9 (F − 32) < 130; −22 < F < 266 −30 < 5 over 9 (F − 32); −22 < F 5 over 9 (F − 32) < 130; F < 266

OpenStudy (danjs):

convert both of those max and min temp values in Celcius, over to F using that given relationship

OpenStudy (anonymous):

OKAY. I am doing that now.

OpenStudy (danjs):

Lower Celcius < FOrmula C to F < upper celcius they want the inequality like that... the celcius relationship

OpenStudy (anonymous):

What will the answer be?

OpenStudy (danjs):

then solve that for F in the middle, or just figure the two limits for F by converting both bounds -30 = (5/9)*(F - 32) 130 = (5/9)*(F - 32)

OpenStudy (danjs):

greater than the first , less than the second

OpenStudy (anonymous):

So that is a?

OpenStudy (danjs):

really dont have to calculate since all the answers give you -22 for lower and 266 for upper F degrees

OpenStudy (danjs):

No, A) says something is less than -30 and larger than 130 at the same time, that is not possible

OpenStudy (anonymous):

okay does that make it B?

OpenStudy (danjs):

yes, the temp range is between those values, a compound inequality like that can be written for it then

OpenStudy (danjs):

that form means both are true at the same time

OpenStudy (anonymous):

Thank you! You are a great teacher.

OpenStudy (danjs):

A says somethign similar to , a number is smaller than 1 and bigger than 1 at the same time, that form requires an OR statment and two seperate inequalities

OpenStudy (danjs):

welcome, anytime